Hello-
I am doing a review on algebra 1 and I am in need of some help.
This is the problem that I need help on:
"Write the equation of line AC, passing through A(1,4) and C (3,6)."
Thank you if you can help!

Answer by venugopalramana(3286)

Hello-
I am doing a review on algebra 1 and I am in need of some help.
This is the problem that I need help on:
"Write the equation of line AC, passing through A(1,4) and C (3,6)."
formula for eqn.of line connecting 2 points...x1,y1...and...x2,y2...is
(y-y1)/(y2-y1)=(x-x1)/(x2-x1)
(y-4)/(6-4)=(x-1)/(3-1)
y-4=x-1
y=x+3